OpenStudy (anonymous):

Lyra and Donna are testing the two-way radios they built for their high school science project. Lyra goes to the top of a building that is 22 meters high with one of the radios and stands directly above the building’s entrance. Donna walks out of the entrance, perpendicular to the building, with the other radio. If Donna is more than 50 meters from the entrance, the connection between the radios breaks. What is the range of communication for the two radios to the nearest meter? 50 meters 55 meters 62 meters 72 meters

OpenStudy (anonymous):

Pythagorean theorem-- c^2 = a^2 + b^2, and if a = 22 and b = 50, then what is c? It's 54.6 If it's plato then round up to 55.

OpenStudy (anonymous):

Thank you so much.. I didn't even think of using that. @Jesusisblack50